Matt is watching a jet take off when he hears the engine roar with an intensity of 100 W/m^2. At what minimum distance should Matt be, in order for him to not hear the jet?

1
Expert's answer
2022-03-24T11:00:20-0400

Explanations & calculations


  • At a distance where the intensity is near his threshold hearing intensity (on average "\small 10^{-12}\,Wm^{-2}" ) he would not hear the roar.

"\qquad\qquad
\begin{aligned}
\small 100\,Wm^{-2}&=\small \frac{Power}{4\pi r^2}\\
\small power&=\small 100(4\pi r^2)\\
\\
\small 10^{-12}\,Wm^{-2}&=\small \frac{power}{4\pi x^2}\\
\small x^2&=\small \frac{100(4\pi r^2)}{4\pi.10^{-12}} \\
\small x &=\small 10^7r
\end{aligned}"

  • He should be at a distance far "\small 10^7" times where he was initially.